### v1

# RemixFusion: Residual-based Mixed Representation for Large-scale Online RGB-D Reconstruction

Abstract.
The introduction of the neural implicit representation has notably propelled the advancement of online dense reconstruction techniques. Compared to traditional explicit representations, such as TSDF, it substantially improves the mapping completeness and memory efficiency. However, the lack of reconstruction details and the time-consuming learning of neural representations hinder the widespread application of neural-based methods to large-scale online reconstruction. We introduce RemixFusion, a novel residual-based mixed representation for scene reconstruction and camera pose estimation dedicated to high-quality and large-scale online RGB-D reconstruction. In particular, we propose a residual-based map representation comprised of an explicit coarse TSDF grid and an implicit neural module that produces residuals representing fine-grained details to be added to the coarse grid. Such mixed representation allows for detail-rich reconstruction with bounded time and memory budget, contrasting with the overly-smoothed results by the purely implicit representations, thus paving the way for high-quality camera tracking. Furthermore, we extend the residual-based representation to handle multi-frame joint pose optimization via bundle adjustment (BA). In contrast to the existing methods, which optimize poses directly, we opt to optimize pose changes. Combined with a novel technique for adaptive gradient amplification, our method attains better optimization convergence and global optimality. Furthermore, we adopt a local moving volume to factorize the whole mixed scene representation with a divide-and-conquer design to facilitate efficient online learning in our residual-based framework. Extensive experiments demonstrate that our method surpasses all state-of-the-art ones, including those based either on explicit or implicit representations, in terms of the accuracy of both mapping and tracking on large-scale scenes. The code will be released at project page.
1. INTRODUCTION
Online dense reconstruction based on RGB-D cameras has made significant advances in recent years and lately, the progress has been propelled by the fast development of robust tracking based on randomized optimization (Zhang et al., 2021; Tang et al., 2023; Zhang et al., 2022) and scalable mapping based on neural map representation (Wang et al., 2021; Chen et al., 2022; Li et al., 2023; Wang et al., 2022; Azinović et al., 2022). The latter has been incurring an in-depth revolution in scene representation methods, from explicit volumetric fields or point clouds to implicit neural fields of occupancy or radiance.
Implicit scene representation leads to improved completeness of dense reconstructions, yet it also introduces new challenges in scene mapping and pose tracking. These challenges are particularly evident in large-scale scene reconstruction. As the size of the scenes scales up, it becomes increasingly difficult to balance the high-precision reconstruction of details with the significant computation and memory overheads required by online systems. In the traditional volumetric fusion approach (Izadi et al., 2011), depth maps are directly fused into a TSDF field. The reconstructed details can be aligned with the depth maps as long as sufficient TSDF resolution is adopted albeit at the expense of higher storage cost. When working with implicit scene representations (Sucar et al., 2021; Zhu et al., 2022; Wang et al., 2023; Johari et al., 2023; Zhang et al., 2023), however, this issue is not straightforward to address as reconstruction details are encoded with neural networks. Even with larger networks and extended training time, encoding high-frequency geometric information remains difficult, and the problem is exacerbated as the scene scales up. Consequently, to make an online dense reconstruction system real-time capable with a limited memory footprint, neither explicit nor implicit scene representation can balance between effective modeling of fine-grained details and efficient reconstruction of scenes at scale.
Lack of details in real-time reconstruction also adversely affects camera tracking in an online reconstruction system, which consequently decreases the overall reconstruction quality. Current methods mostly rely on optimizing a rendering loss in a frame-to-model approach for camera pose estimation (Zhu et al., 2022; Wang et al., 2023). In large-scale reconstructions, excessively flat optimization gradients lead to poor convergence. This issue is particularly prominent in multi-frame pose optimization based on bundle adjustment, causing severe optimization oscillations and failure to achieve the desired joint optimization effect, as we will show in the experiments (see Table 9 and Figure 12).
Residual learning has been proven in various works (He et al., 2016; Xiangli et al., 2022; Kim et al., 2016) to improve the convergence of neural network training, enhance generality, and enable the network to encode rich details. These advantages are especially useful when learning implicit representations for large-scale scene reconstruction. Motivated by that, we propose RemixFusion, adopting the concept of residuals with a mixed representation of implicit and explicit, for both detail-rich mapping and accurate pose tracking in online dense reconstruction. To address the inefficiency of implicit representations in capturing high-frequency details in large scenes, we first combine explicit and implicit scene representations in a residual form. The base of our mixed representation is a coarse-grained explicit 3D grid volume that stores low-frequency scene structure. High-frequency geometry details are captured in the parameters of the neural module, coupling with the coarse-grained base as a residual. This approach reduces memory overhead by lowering the resolution of the explicit representation while preserving scene details in a memory-efficient form through neural representation. Furthermore, the training complexity of the residual neural module is significantly reduced since it focuses on encoding only high-frequency features. As a result, this mixed representation not only preserves fine-grained details in a memory-efficient way but also greatly improves online learning efficiency.
For pose estimation, based on the mixed representation, we propose a residual-based multi-frame pose optimization method. Specifically, we estimate residual pose corrections to refine the initial poses from front-end tracking, thereby enhancing geometric consistency during bundle adjustment (BA). By encoding only residual poses in the network, our approach allows the BA based on implicit representations to focus on optimizing pose changes rather than absolute poses, thus improving the learning efficiency and multi-view consistency. In practice, the real-time constraints make it impractical to perform bundle adjustment on the whole reconstruction. The sparse sampling of the scene results in discontinuous perception of surface details during joint optimization, which in turn reduces the optimization’s ability to escape local minima. To address this challenge, we introduce an adaptive gradient amplification based on the reconstructed surface, allowing the BA to obtain better optimization gradients even when the detail perception is discontinuous in real-world large-scale scenarios.
As shown in Figure 1, our online dense reconstruction method, employing an effective fusion of explicit and implicit representations, achieves fine-grained reconstruction of large-scale scenes with a relatively low GPU memory cost. Our method exhibits significantly improved online learning efficiency thanks to the explicit map, resulting in a frame rate that is 2.5 times higher than that of other methods based solely on implicit representation. Furthermore, the residual-based camera pose bundle adjustment benefits pose tracking, which improves tracking accuracy by over the state-of-the-art dense SLAM (Johari et al., 2023) on BS3D (Mustaniemi et al., 2023). In summary, our contributions include:
We propose a mixed residual-based representation for dense RGB-D reconstruction of large-scale scenes, which preserves fine-grained details with relatively low memory and computational cost.
We propose a residual-based bundle adjustment technique that employs a tiny MLP for residual-based pose refinement. Compared to traditional BAs, our method improves pose estimation in terms of both efficiency and robustness.
We have implemented an efficient system of online RGB-D dense reconstruction which realizes robust and fine-grained real-time reconstruction for large scenes over 1000 with an affordable GPU memory footprint.
2. RELATED WORK
The field of online reconstruction methods constitutes a substantial area of research. In this context, we review the most relevant literature on large-scale indoor scene reconstruction, including representations based on both explicit and implicit neural methods.
Explicit methods.
The explicit scene representation has been widely studied in the last decades, most including voxel-based and point-based methods. In the context of voxel-based methods, KinectFusion (Izadi et al., 2011) proposes the Truncated Signed Distance Function (TSDF) for encoding the scenes with a consumer depth sensor. Considering the waste of encoding the empty space of the scene, follow-up works (Whelan et al., 2012; Roth and Vona, 2012) extend the KinectFusion by leveraging a moving TSDF volume to encode larger scenes. (Nießner et al., 2013a; Dai et al., 2017b) encode the scenes with hash blocks to further improve the scalability of scene reconstruction. Global bundle adjustment is utilized to reduce the drift in pose estimation and enhance model consistency in terms of large-scale indoor environments (Dai et al., 2017c). For points-based methods, (Keller et al., 2013; Whelan et al., 2015, 2016) leverages points and surfels to achieve scalability and flexibility for encoding scenes. With explicit scene representation, these methods leverage classic second-order optimization methods, such as Gaussian-Newton or Levenberg-Marquardt (LM). Recently, (Zhang et al., 2021, 2022) propose random optimization to improve the robustness of camera tracking and demonstrate good performance on large-scale fast-moving motions. However, explicit methods suffer from memory consumption, especially in large-scale scenes.
Implicit methods.
Inspired by NeRF (Mildenhall et al., 2021), the pioneering work of NeRF-based SLAM, iMAP (Sucar et al., 2021) proposes to use keyframe-based joint optimization via differential volume rendering and encode the entire scene in a multilayer perceptron (MLP). To improve geometric details, NICE-SLAM (Zhu et al., 2022) and Vox-Fusion (Yang et al., 2022) incorporate multi-scale feature grids with shallow decoders to collaborate on memorizing the geometry and texture. Recent advanced works, Co-SLAM (Wang et al., 2023), ESLAM (Johari et al., 2023) and Point-SLAM (Sandström et al., 2023) exploit more efficient parametric embedding like multi-resolution hash encoding (Müller et al., 2022), multiscale feature planes (Chan et al., 2022) and data-driven neural point clouds (Xu et al., 2022b) for memory efficiency and faster convergence speed. These approaches seamlessly integrate pose estimation and reconstruction by a render-and-align pattern. (Xu et al., 2022a) addresses robust pose estimation of implicit representations of Hermite Radial Basis Functions (HRBFs). Alternative approaches decouple tracking from reconstruction and seek more accurate and robust pose estimation, paying less attention to the reconstruction quality. Feature matching is utilized in traditional approaches (Chung et al., 2022), while neural approaches (Koestler et al., 2022; Teed and Deng, 2021; Zhang et al., 2023) leverage the deep multi-view stereo or learnable optical flow estimator to achieve frame-to-frame registration. However, these methods struggle to preserve real-time and detailed reconstruction in large scenes.
Representations in large scenes.
Reconstruction of large-scale scenes requires efficient representations. As for explicit voxel-based methods, hashing schemes (Nießner et al., 2013b; Dai et al., 2017c) and octrees (Liu et al., 2020; Mao et al., 2023) are adopted to reduce the memory footprint. However, these methods do not reduce the number of parameters in modeling and lack the ability of completion. In terms of implicit methods, learnable hash grids (Wang et al., 2023; Zhang et al., 2023), tri-planes (Johari et al., 2023) and neural points (Sandström et al., 2023; Hu et al., 2023) are utilized to achieve the trade-off of efficiency and reconstruction quality. However, these methods suffer from the high latency of the model update. Another line of work leverages submaps (Tancik et al., 2022; Tang et al., 2023; Mao et al., 2023) to dynamically allocate the implicit maps with the moving camera. The multi-submap techniques improve the scalability while costing much more time in submap management. Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) shows promising results in novel view synthesis. A series of works (Keetha et al., 2024; Huang et al., 2024a; Matsuki et al., 2024; Ha et al., 2024) explore applying 3DGS to monocular or RGB-D SLAM. Submaps and loop closure techniques (Zhu et al., 2025) are incorporated to improve the tracking accuracy. Furthermore, a more compact representation (Peng et al., 2024) is proposed to enhance the efficiency of 3DGS. These methods demonstrate high-fidelity rendering and generalization capabilities in different scenarios. However, these 3DGS-based methods are still less memory-friendly than implicit methods since they need numerous 3DGS for the explicit model, especially in large-scale scenes. We propose to take advantage of both explicit and implicit representations to remain both memory-efficient and time-efficient via residual mixing.
3. METHOD
3.1. Overview
The proposed modules in RemixFusion are necessary and beneficial to each other.
In this section, we first introduce our residual-based framework, including the explicit-implicit mixed scene representation and residual-based bundle adjustment for camera poses. Our insight behind these designs is that the residual can be adopted to force networks to focus on the high-frequency details upon existing coarse geometry, relieving the burden of neural learning and enhancing efficiency. In this way, the proposed online dense reconstruction framework is both storage-efficient and time-efficient, which is crucial for application in large-scale scenes with real-time requirements.
More specifically, as shown in Eq. 1, we decouple the complete representation into a coarse representation and a residual one , which are optimized during reconstruction and pose estimation. In general, is expected to be fast and robust, such as TSDF, providing the coarse geometry seamlessly with the input data. is supposed to be expressive and capture the fine-grained details based on . The coarse estimation is usually constructed based on explicit representations, and the neural networks are adopted for residual refinement . The coarse estimation and residual refinement are performed alternately for both reconstruction and pose estimation.
where denotes operation to aggregate and .
The overview of our method is demonstrated in Figure 2. Given the successive RGB-D inputs, we first leverage the scalable randomized optimization with limited and fixed footprints for the coarse pose estimation (Figure 2(a)). Keyframes are selected from these input frames. The corresponding initial camera poses are fed into a tiny MLP for the joint residual pose refinement (Figure 2(b)).
Note that the proposed residual-based BA of the camera poses is optimized by the gradients given by the multi-view constraints based on the reconstructed scene. To support pose estimation better, an explicit-implicit scene representation mixture is introduced simultaneously. The online learning (Figure 2(c)) of this representation depends on the optimized camera poses. Specifically, we perform ray casting and sample many points along the ray directions accordingly. Embeddings of these points, including the hash and explicit embedding together with positional embedding, are sent to two MLPs for the TSDF and RGB residual prediction. The residuals indicating the high-frequency details are added to the explicit coarse bases, outputting the final reconstruction. The residual-based BA and residual-based mapping are running alternately with continuously updated global reconstruction and optimized keyframe poses.
In the following sections, we first introduce the residual-based mapping (Section 3.2), and then demonstrate the optimization of residual-based BA based on the global residual-based map (Section 3.3). The strategies of the scalable randomized optimization and other implementation techniques are illustrated in Section 3.4.
3.2. Residual-based Mapping
In large-scale scenes, various layouts and complex structures make it challenging for the scene representations to remain both expressive and efficient. Shown in Figure 2(c), we employ a mixed representation to address this challenge. This representation consists of a explicit coarse TSDF grid and an implicit neural map , which correspond to and in Eq. 1 respectively. The key insight of the residual-based representations is to obtain the coarse reconstruction in real time and provide the basis surface for the implicit module to efficiently deform and refine. Consequently, given an arbitrary 3D position , the attributes which are the geometry and appearance of can be formed in a residual-based aggregation as:
where denotes the trilinear interpolation operation to query RGB and TSDF values for and is the decoder for the implicit neural map .
Mixed Scene Representation
Given the posed RGB-D frames, the explicit component of the proposed scene representation is constructed via TSDF Fusion (Curless and Levoy, 1996) with a relatively low resolution to store the coarse TSDF and RGB attributes, serving as explicit bases of our reconstruction. Similar to some previous implicit-based methods, the implicit component includes two kinds of embedding functions, which are hash embedding (Müller et al., 2022) and positional embedding (Müller et al., 2019), to encode the reconstruction residuals for an arbitrary position upon the . Specifically, adopting this joint embedding in the representation can balance the training efficiency and reconstruction quality(Wang et al., 2023), which is critical for online reconstruction. Furthermore, the residual learning can not be independent without the explicit bases , the final residual-based aggregation to calculate the attributes for a position can be formed based on Eq. 2 as:
where denotes the trilinear interpolation operation for on the grid to query the coarse RGB and TSDF values. denotes the positional embedding.
Neural Learning of the implicit residual
Different from the can be reconstructed directly based on the fusion of the posed RGB-D frames, and in our implicit neural map need to be carefully optimized during the reconstruction ( is a constant projection). Therefore, proper supervision for learning and formulating a good loss function is critical here.
Following (Sucar et al., 2021; Zhu et al., 2022; Wang et al., 2023), we adopt the input RGB-D frame sequence with the estimated poses as the approximation of ground-truth modeling to provide the supervision in a projection-and-measure fashion. Volume rendering is leveraged to output the re-projected RGB-D images based on the weights of points on the casting ray from each image pixel. Similar to (Azinović et al., 2022), the weights are obtained by the Sigmoid functions and as shown in Eq. 5.
Given an image pixel and the corresponding sampled points on its casting rays, the rendered RGB values and depth values can be obtained as Eq. 6 and Eq. 7.
where is the 3D position of under the camera coordinates with the corresponding pose and denotes truncation threshold adopted in the TSDF construction.
One more thing, due to the resolution of being low, the truncation threshold for it needs to be larger than the one adopted in the residual component . To align these two modules, the explicit bases adopted in TSDF component of Eq. 3 needs to be specified as:
The is the clamp function which ensures to be in the range of with threshold . The aligned explicit bases is only adopted for the TSDF calculation, while with the direct trilinear interpolation is proper for the RGB prediction.
Figure 3. Fully described in the text.
For the running efficiency, we sample pixels in each iteration for neural learning. More details about sampling strategies can be found in Section 3.4. We use the observed RGB and depth images for supervision. The photometric and geometric loss are as follows:
Following (Wang et al., 2023; Johari et al., 2023), we leverage an approximate TSDF loss for points within the truncation () and a free-space loss for points in front of the surface (), where refers to the sampled depth values on the rays. These two designs are aimed at recovering the accurate and detailed geometry using depth observations as approximate ground-truth TSDF supervision.
where is the approximate ground-truth TSDF value for the pixel , and () denotes the sampled points allocated within and farther than a pre-defined distance threshold to the observed approximate surface .
Besides these loss functions, a smoothness loss is applied to ease the hash collisions, avoiding the noisy points in empty space and make the predictions smooth.
where is the random sampled vertices on the hash grids and means the difference of hash features between the adjacent vertices along these three axes.
In general, the mixed mapping is carried out with iterations every frames, and the loss function under observation and its corresponding poses is defined as Eq. 13. are the corresponding weights for each loss component.
Figure 3 illustrates that the mixed representations focus on the high-frequency details and completion based on the coarse representations. Moreover, Figure 4 illustrates the better efficiency of RemixFusion in reconstruction. We, in essence, achieve a better trade-off by the proposed residual-based representation, which takes advantage of both explicit and implicit representations.
RemixFusion can reconstruct the scene faster than Co-SLAM, demonstrating the efficiency of the proposed representations.
3.3. Residual-based Bundle Adjustment
As mentioned above, the residual-based mapping is based on the estimated camera poses, which are given by online camera tracking. For the best efficiency, the pose estimation can be obtained through an existing randomized optimization-based method (Zhang et al., 2021) for each RGB-D frame. Different from the gradient-based optimization used in (Bylow et al., 2013), the randomized optimization is more robust even if the optimization is of high nonlinearity. This method aims to minimize the distance between the transformed frame depth based on the estimated pose and the zero-crossing surface geometry of the reconstruction, which forms a frame-to-model optimization. Details of this method can be found in the supplementary materials.
However, there is still inevitable accumulated drift in front-end estimated camera poses, which addresses the necessity of multi-view bundle adjustment (BA), especially for large-scale scenes.
Bundle adjustment for residual pose
In contrast to the reconstruction, the bundle adjustment optimizes the camera poses with the neural implicit map fixed. Generally, the objective loss functions are the same as Eq. 13 based on obtained by Eq. 3 for sampled points . The difference is that we further optimize the camera poses for the observed RGB-D frames while the mixed mapping is fixed. The goal of bundle adjustment is formulated as:
Bundle adjustment (BA) in previous neural SLAM and traditional alternatives are trying to optimize for each frame directly based on the back-propagation of the gradients given by Eq. 14, which makes the optimization for each frame relatively independent. Lack of global awareness among independent optimization of for each frame may lead to conflicts under the multi-view consistency constraints. These conflicts would limit the magnitude of optimization in BA to promise consistency. Therefore, we propose to adopt a single MLP encoding individual poses to be optimized for better global awareness in BA.
However, using the MLP to encode the complete camera poses only based on the frame index is inefficient and difficult for training. Inspired by the residual-based mixed representation (Section 3.2), we propose to change the output of from the complete pose to only the residual pose, aiming to reduce network learning complexity and thereby enhance training efficiency. Specifically, we encode the pose changes along with the frame index in . Note that the frame index should be normalized to according to the number of keyframes for better convergence. The residuals are added back to the initial poses to gain the final results . Compared to the independent optimization of each keyframe, the MLP is more comprehensive, and the residual-based BA is more globally aware and efficient. The process of the residual-based BA can be formulated as Eq. 16. The and correspond to and in Eq. 1 respectively. The residual-based BA is running with iterations every frames. denotes the normalization of the frame index for better convergence.
Adaptive gradient amplification of the BA module
Although the residual-based BA is global-aware and efficient, there is still room for improvement to BA in large-scale scenes. Due to the real-time and memory requirements, the sampling has to be sparse. Meanwhile, multi-view constraints of BA are applied only to a subset of all pixels, which makes the optimization of Eq. 14 easy to get stuck in the local optima, especially in large-scale scenes.
Inspired by the Simulated Annealing algorithm (SA), gradient amplification (GA) seems like a good choice to overcome this problem. The key idea is to amplify the optimization gradients near the reconstructed surface and encourage the BA optimization to explore more. The simplest way to amplify the gradients is to move the cameras in the direction they are facing or in a random direction, by a certain distance. This would lead to misalignment of the currently observed surfaces and reconstructed surfaces. The lack of consideration of the 3D geometric structures causes the BA module to fail to achieve the expected results due to gradient amplification disorder.
We propose an adaptive gradient amplification technique based on the reconstruction geometry, which amplifies the optimization gradients of the BA process near the zero-crossing surface, further improving the accuracy of pose estimation. Specifically, we change Eq. 8 to Eq. 17 with multiplied by only in BA , which results in the imbalance of the predicted and approximate ground-truth TSDF values, which should both be equal to 1 in the free space. In this way, the predicted TSDF absolute values are clamped with the threshold (¿1). According to Eq. 11, the loss and are amplified to drive the observed surfaces to get close to the zero-crossing surface. The residual-based BA would then try to minimize the amplified losses and achieve a more multi-view consistent balance.
The proposed optimization momentum improves the performance of bundle adjustment, which is validated by the aligned reconstruction.
Figure 5(a) visualizes the optimization process of the bundle adjustment on the scene foobar of BS3D. The states, colorized by the errors of pose estimation, are 6D poses of all keyframes, which are visualized in 2D. The optimization starts with the poses of front-end tracking. The red and blue regions indicate high and low errors, respectively. BA with the proposed GA converges fast to the global minima, while BA without GA tends to be stuck in the local minima, and the improvement is marginal. The visualization is shown by the alignment of point clouds generated with the optimized poses and ground-truth poses, which correspond to the yellow and blue ones in Figure 5(b) . Overall, the proposed GA helps the bundle adjustment in large-scale scenes explore more and converge to the global minima. Extensive experiments show that the residual-based BA is efficient and robust, which can steadily refine the estimated initial camera poses with more global comprehension.
3.4. System Implementation
We illustrate the proposed scalable pose estimation for large-scale scenes using randomized optimization and system implementation, including sampling and keyframe selection, in this section.
Scalable Randomized Optimization
First, we review the core insight of randomized optimization used in (Zhang et al., 2021), as well as the basic principles and cost functions of pose estimation. In brief, the goal is to provide accurate 6DoF camera pose for live RGB-D frames . We first back-project the pixel with depth to its 3D position in the camera space, its point-to-surface distance in the world space could then be queried using . defines the difference between the queried TSDF values from the moving volume and the approximated ground-truth TSDF supervision, which is similar to Eq. 10 and Eq. 11. The minimization of these point-to-surface distances described in Eq. 18 is tackled by the randomized optimization with iterations using a pre-sampled particle swarm template (PST). We encourage readers to refer to supplementary materials for more details.
where denotes the depth image and means the query of the reconstructed TSDF volume.
One main difference of our method against ROSEFusion (Zhang et al., 2021) is that we leverage a moving volume to perform pose estimation to make our tracking module scalable to large-scale scenes. This is inspired by the success of the moving TSDF-Fusion strategy (Whelan et al., 2012; Roth and Vona, 2012). The process of the management of the moving explicit volume and pose estimation based on randomized optimization is illustrated in Algorithm 1. We initialize the first moving volume given the first camera position , and the input RGB-D frames , which are observed within volume are integrated. The camera pose would be marked as an anchor each time the volume is moved. We transform the moving volume only when the Euclidean distance between the current camera and the last anchor camera exceeds a threshold in any direction of the 3 axes. We create a new moving volume , inheriting the overlapping regions from to minimize the transferring costs, shown in Eq. 19. After swapping, is abandoned.
where refers to the voxel coordinate and is the default value of the TSDF volume. Here we set its TSDF value to one and the other attributes, including color and weights, to zero. In our implementation, we use a dense axis-aligned moving TSDF volume with resolution higher than that of . Our system can accurately perceive local environments with a fixed and moderate memory cost to maintain efficiency.
Sampling and Keyframe Selection
Following Co-SLAM (Wang et al., 2023), we downsample each keyframe and only save 5 of the pixels in the keyframe database that contains the RGB, depth and pose data. The mapping process is continuously running every frame, and we sample pixels from the keyframe database each time. After sampling, we can back-project the pixels to 3D space, and sample points along the rays of different pixels by ray casting, which contains points uniformly sampled and sampled near the observed surface. Then, we can get the points , which are points in total to be queried and learned.
4. EXPERIMENTS
4.1. Experimental Setup
Dataset
We evaluate our method on two large-scale indoor datasets BS3D (Mustaniemi et al., 2023) and uHumans2 (Rosinol et al., 2020). We also provide the comparison for the room-level publicly available dataset: Replica (Straub et al., 2019), TUM RGB-D (Sturm et al., 2012), ScanNet (Dai et al., 2017a), and FastCaMo (Zhang et al., 2021). This evaluation aims to benchmark our tracking and reconstruction performance against current state-of-the-art methods. BS3D is a challenging real-world large-scale RGB-D dataset, comprising over 10,000 RGB-D images annotated with ground-truth trajectories from a motion capture system. uHumans2 is a large-scale publicly available synthetic dataset with multiple rooms and complicated layouts, presenting a significant challenge for SLAM. We evaluate eight large-scale sequences of BS3D, two sequences of uHumans2, alongside commonly referenced sequences from Replica, ScanNet, TUM RGB-D, and FastCaMo-Synth (noise-free). Furthermore, additional qualitative comparisons are conducted on BS3D, uHumans2, FastCaMo-Large proposed in MIPS-Fusion (Tang et al., 2023), and self-captured sequences to highlight the different performance on large-scale scenes.
Metrics
In terms of 3D reconstruction quality, we follow (Zhu et al., 2022) and adopt Accuracy(cm), Completion(cm), and Completion ratio(%) with the threshold of 5cm and 10cm as well as F1-score. Following (Azinović et al., 2022; Zhu et al., 2022), we filter the noisy points that are not visible in any observation with ground-truth depth images for a fair comparison, as many neural approaches tend to predict numerous noisy points in empty spaces. Then we use the Iterative Closest Point (ICP) for alignment with ground-truth meshes. Metrics like PSNR, SSIM, LPIPS, and Depth-L1 (m) for 2D rendering comparison are adopted following (Keetha et al., 2024). As for camera tracking, the ATE RMSE (Sturm et al., 2012) is adopted. The system FPS (frames per second) and GPU memory usage are considered for efficiency comparison.
Methods | cafeteria | corridor | foobar | hub | juice | lounge | study | waiting | Avg. | |
Implicit | iMAP∗ | – | 255.4 | 295.8 | 184.6 | 64.7 | 299.0 | 318.5 | 143.6 | 262.1 |
NICE-SLAM | 52.9 | 8.7 | 13.1 | 12.1 | 4.9 | 11.0 | 7.5 | 6.6 | 14.6 | |
Co-SLAM | 127.6 | 11.5 | 106.5 | 6.2 | 5.2 | 48.8 | 6.5 | 19.8 | 41.5 | |
MIPS-Fusion | 122.1 | 78.5 | 41.2 | 36.5 | 6.2 | 11.3 | 6.1 | 19.7 | 40.2 | |
ESLAM | 7.7 | 5.7 | 10.4 | 4.3 | 4.1 | 6.3 | 4.7 | 8.3 | 6.4 | |
Explicit | ElasticFusion | 193.1 | 230.6 | 64.2 | 76.9 | 119.3 | 314.0 | 85.5 | 142.4 | 153.3 |
BundleFusion | – | – | – | – | 10.0 | – | – | – | – | |
BAD-SLAM | – | 170.0 | 334.1 | – | 22.4 | 8.3 | 4.7 | – | – | |
SplaTAM | – | – | 168.1 | 260.3 | 31.2 | – | 372.5 | 63.5 | – | |
MonoGS | 448.4 | – | 491.2 | 207.1 | 182.8 | 317.2 | 211.0 | 211.4 | – | |
LoopSplat | – | – | 31.5 | – | – | 149.1 | – | – | – | |
Photo-SLAM | 15.4 | 23.2 | 26.6 | 7.6 | 4.0 | 24.6 | 6.0 | 6.5 | 14.3 | |
RTG-SLAM | 21.9 | 11.6 | 11.1 | 8.4 | 9.9 | 12.0 | 6.7 | 15.6 | 12.2 | |
GS-ICP SLAM | 35.1 | 23.9 | 19.9 | 9.2 | 6.3 | 24.0 | 5.8 | 7.3 | 16.4 | |
ORB-SLAM3 | 5.6 | 6.7 | 7.5 | 6.1 | 6.0 | 15.8 | 5.4 | 2.9 | 7.0 | |
RemixFusion | 6.8 | 6.3 | 5.8 | 4.5 | 3.1 | 4.2 | 3.2 | 3.0 | 4.6 | |
RemixFusion-lite | 10.1 | 11.0 | 6.8 | 5.0 | 4.1 | 4.6 | 4.3 | 5.1 | 6.4 |
Baselines
We compare our method to both the implicit and explicit state-of-the-art methods. The former includes NeRF-based methods like iMAP (Sucar et al., 2021), NICE-SLAM (Zhu et al., 2022), Co-SLAM (Wang et al., 2023), ESLAM (Johari et al., 2023) and MIPS-Fusion (Tang et al., 2023). The latter indicates ElasticFusion (Whelan et al., 2015), BundleFusion (Dai et al., 2017c) and BAD-SLAM (Schops et al., 2019). RTG-SLAM (Peng et al., 2024), GS-ICP SLAM (Ha et al., 2024), SplaTAM (Keetha et al., 2024), MonoGS (Matsuki et al., 2024), Photo-SLAM (Huang et al., 2024a) and LoopSplat (Zhu et al., 2025) are encompassed as the representative of the SOTA 3DGS-based SLAM approaches. For more comprehensive comparisons, the sparse SLAM framework ORB-SLAM3 (Campos et al., 2021) and DROID-SLAM (Teed and Deng, 2021) that specialize in robust camera tracking are included. The implementation of iMAP* is from NICE-SLAM. ESLAM, SplaTAM, and MonoGS are evaluated using images at half resolution as inputs on BS3D and uHumans2 datasets due to excessive GPU occupancy. For RTG-SLAM, a modified version was employed due to its high GPU memory allocation in global optimization on these two datasets. More details can be found in the supplementary materials.
Implementation Details
We evaluate RemixFusion on a desktop PC equipped with a 3.90GHz Intel Core i9-14900K CPU and an NVIDIA RTX 3090 Ti GPU. In terms of camera tracking, we use iterations per frame for the randomized optimization. We implement the scalable randomized optimization based on the PyCUDA (Klöckner et al., 2012) libraries for acceleration. The tracking process is parallel to mapping. For Mapping, we sample pixels and points along each ray, including 11 for uniform samples and 48 for samples near the surface. The resolution is 200 for and for in BS3D. The threshold used in residual-based BA is and . The mapping and BA process are iteratively performed with iterations. More details can be found in the supplementary materials.
4.2. Quantitative and Qualitative Comparison
In this section, we present quantitative and qualitative results, including the camera pose estimation, 3D mesh reconstruction, and 2D rendering on different datasets. The lightweight version of our method, requiring fewer optimization iterations for both tracking and mapping, is denoted as RemixFusion-lite. Details can be found in the supplementary materials.
Methods | Metrics | FPS | ||
Acc. | Comp. | Comp. Ratio(%) | ||
Co-SLAM | 4.57 | 3.91 | 85.30 | 20 |
MIPS-Fusion | 6.93 | 16.78 | 60.15 | 9 |
ESLAM | 5.92 | 8.40 | 85.73 | 5 |
RTG-SLAM | 6.71 | 8.28 | 63.68 | 5 |
GS-ICP SLAM | 9.09 | 8.37 | 53.44 | 26 |
RemixFusion | 4.34 | 3.56 | 86.88 | 27 |
RemixFusion-lite | 4.70 | 4.06 | 83.45 | 94 |
4.2.1. Real-time Reconstruction Regardless of Tracking.
Table 2 presents the results of 3D reconstruction using ground-truth poses to eliminate the tracking effects in SLAM. Thanks to the residual-based mixed representation, our method is the most competitive one in accuracy (4.34cm) and completeness (3.56cm), surpassing the state-of-the-art by 5 in accuracy and 9 in completeness. The second-best method, after our own, is Co-SLAM, which achieves the accuracy of 4.57cm and the completeness of 3.91cm. While the completeness ratio of ESLAM is the second-best for implicit methods, its completeness and accuracy remain suboptimal, indicating the incompact reconstruction. Notably, with ground-truth poses, the best 3DGS-based method demonstrates inferior reconstruction performance (31.9 worse than the best implicit method in accuracy) with the online setting. RTG-SLAM leverages an efficient and compact 3DGS representation to save memory, but the reconstruction performance is not desirable, indicating the trade-off between efficiency and accuracy. Although GS-ICP SLAM has a comparable mapping FPS to our method, its reconstruction accuracy is 4.75 worse than ours. This indicates that achieving high-quality geometric reconstruction with 3DGS requires significantly more optimization iterations in real-world large-scale scenarios, which may lead to compromised 3D reconstruction under real-time SLAM constraints. Note that the outputs of MIPS-Fusion are obtained by running the official code, and some objects are missing. This leads to worse performance of completeness and comparable performance of completeness ratio. The lightweight version of our method only performs mapping with a few optimization iterations in this experiment and is significantly faster, with only a slight decrease in reconstruction accuracy (-0.36cm) and completeness (-0.5cm), which further proves the effectiveness and robustness of our method.
In Figure 6, we present the visualization of the reconstruction of different approaches. There are many artifacts for RTG-SLAM and GS-ICP SLAM, primarily due to the discontinuity in rendered depths. Our method excels in preserving detailed geometry, such as the leaves on the floor (the second column) and the pillows (the fourth column). Additionally, the comparison of ours and Co-SLAM (the best approach except for ours in Table 2) with camera poses of RemixFusion can be found in the supplementary materials. The improvement compared to the implicit methods proves that the residual-based mixed representation enhances the reconstruction quality with limited time, which is crucial for SLAM.
4.2.2. Pose Estimation.
Table 1 shows ATE RMSE (cm) on 8 sequences of BS3D for our method compared to five state-of-the-art implicit RGB-D SLAM and three explicit traditional methods. Evaluations of the state-of-the-art 3DGS-based methods (SplaTAM, MonoGS, RTG-SLAM, GS-ICP SLAM, LoopSplat and Photo-SLAM) and the sparse SLAM (ORB-SLAM3) are also incorporated for a more comprehensive comparison for tracking accuracy.
Our method attains the best tracking performance (4.6cm) for 8 challenging sequences on average while running in real time. There is 28.1 improvement for our method in tracking on average compared to the SOTA methods. Implicit methods struggle with pose estimation in these large-scale scenes. Specifically, implicit methods, except for NICE-SLAM and ESLAM , all fail in tracking on the scene cafeteria, which is more than 400 and the layouts are complex, even including some textureless areas. These factors pose challenges for robust camera pose estimation and real-time reconstruction, resulting in unsatisfactory reconstruction quality, unstable pose estimation, and inadequate bundle adjustment for most methods. Although our method is inferior to the sparse SLAM ORB-SLAM3 on this challenging sequence, our method surpassed ESLAM by 11.7. For the scene hub, our method underperforms the best method by a margin of 0.2cm. Although NICE-SLAM and ESLAM demonstrate stable tracking performance on BS3D, they suffer from low FPS and high GPU memory consumption in large-scale scenes, as shown in Table 8. In contrast, our method maintains accuracy and robustness on these large-scale scenes RemixFusion-lite (over 25FPS), the lightweight version, is the second-best alternative on average, further proving the efficiency and robustness of our method.
Methods | office | apartment | Avg. | ||||||||||||
ATE RMSE | PSNR | SSIM | LPIPS | D-L1 | ATE RMSE | PSNR | SSIM | LPIPS | D-L1 | ATE RMSE | PSNR | SSIM | LPIPS | D-L1 | |
DROID-SLAM | 6.84 | – | – | – | – | 3.98 | – | – | – | – | 5.41 | – | – | – | – |
NICE-SLAM | 14.49 | 21.8 | 0.929 | 0.395 | 0.134 | 6.79 | 25.35 | 0.976 | 0.345 | 0.086 | 10.64 | 23.575 | 0.953 | 0.370 | 0.110 |
Co-SLAM | 1268.23 | 17.79 | 0.838 | 0.752 | 1.885 | 26.08 | 0.977 | 0.412 | 0.171 | 10.95 | 639.59 | 21.94 | 0.908 | 0.582 | 1.028 |
MIPS-Fusion | 56.46 | 19.99 | 0.888 | 0.586 | 0.989 | 16.07 | 22.29 | 0.953 | 0.510 | 0.669 | 36.27 | 21.14 | 0.921 | 0.548 | 0.829 |
ESLAM | 6.53 | 24.32 | 0.948 | 0.198 | 0.129 | 4.13 | 27.91 | 0.984 | 0.170 | 0.065 | 5.33 | 26.12 | 0.966 | 0.184 | 0.097 |
ElasticFusion | 1314.37 | – | – | – | – | 116.42 | – | – | – | – | 715.40 | – | – | – | – |
BAD-SLAM | – | – | – | – | – | 19.54 | – | – | – | – | – | – | – | – | – |
SplaTAM | 1930.69 | 10.29 | 0.264 | 0.719 | 1.564 | 87.32 | 22.35 | 0.766 | 0.384 | 0.221 | 1009.01 | 16.32 | 0.515 | 0.552 | 0.893 |
MonoGS | 26.21 | 20.63 | 0.653 | 0.534 | 1.939 | 6.92 | 32.40 | 0.927 | 0.092 | 0.270 | 16.57 | 26.52 | 0.790 | 0.313 | 1.105 |
Photo-SLAM | – | – | – | – | – | 11.12 | 23.58 | 0.961 | 0.364 | 1.371 | – | – | – | – | – |
LoopSplat | 160.47 | 21.99 | 0.925 | 0.458 | 1.177 | 56.09 | 24.99 | 0.969 | 0.384 | 0.867 | 108.28 | 23.49 | 0.947 | 0.421 | 1.022 |
RTG-SLAM | 330.49 | 16.24 | 0.768 | 0.618 | 2.405 | 90.23 | 20.01 | 0.922 | 0.542 | 1.299 | 210.36 | 18.13 | 0.845 | 0.580 | 1.852 |
GS-ICP SLAM | 4799.29 | 18.84 | 0.612 | 0.585 | 3.132 | 185.17 | 23.33 | 0.784 | 0.411 | 2.571 | 2492.23 | 21.09 | 0.698 | 0.498 | 2.852 |
RemixFusion | 5.44 | 24.66 | 0.951 | 0.368 | 0.092 | 3.93 | 27.46 | 0.983 | 0.326 | 0.077 | 4.69 | 26.06 | 0.967 | 0.347 | 0.085 |
Explicit dense RGB-D methods struggle with robustness in these challenging scenes. Among them, only ElasticFusion, RTG-SLAM GS-ICP SLAM, Photo-SLAM, and ORB-SLAM3 consistently succeed in tracking on all scenes, while BundleFusion and BAD-SLAM encounter failures in certain cases, primarily due to front-end odometry issues, which is also validated in (Mustaniemi et al., 2023). Although ElasticFusion can finish tracking on these sequences, its trajectory proves inaccurate.
3DGS-based methods, including RTG-SLAM, GS-ICP SLAM, and Photo-SLAM, are comparable to NICE-SLAM in tracking. These methods leverage either multi-level ICP or ORB features (Mur-Artal and Tardós, 2017) for tracking, which is decoupled from 3DGS optimization. In contrast, other 3DGS-based methods, including SplaTAM, MonoGS, and LoopSplat, which leverage the rendering loss as the key objective function for camera pose optimization, exhibit unstable tracking performance on this large-scale dataset. Although they can attain high-fidelity rendering given good camera poses, it is challenging to optimize the 3DGS and camera poses simultaneously in large-scale scenes, since there are drastically more 3DGS to optimize. Notably, LoopSplat achieves successful tracking on the foobar and lounge sequences but fails on the remaining sequences on BS3D. SplaTAM fails on 3 sequences and MonoGS fails on 1 sequence. In contrast, our method is robust and surpasses the best 3DGS-based method (RTG-SLAM) by 62.3 for tracking.
ORB-SLAM3, known for sparse SLAM, achieves robust and accurate pose estimation regardless of dense reconstruction. In comparison, our method achieves superior tracking accuracy across most scenes and emerges as the most accurate dense RGB-D SLAM system overall for tracking, surpassing even sparse SLAM systems like ORB-SLAM3 by 34.3. Furthermore, while ORB-SLAM3 focuses on sparse reconstruction, our method excels in dense reconstruction with more accurate pose estimation.
The details, including the leaves of plants and pillows on the sofa, are more fine-grained for RemixFusion.
Methods | Metrics | cafeteria | corridor | foobar | hub | juice | lounge | study | waiting | Avg. |
NICE-SLAM | Acc. | 45.65 | 20.78 | 24.81 | 21.68 | 10.27 | 23.35 | 20.11 | 18.97 | 23.20 |
Comp. | 23.82 | 6.36 | 8.42 | 6.01 | 3.95 | 9.49 | 6.00 | 5.16 | 8.65 | |
Comp. Ratio(%) | 27.84 | 87.09 | 71.61 | 87.48 | 96.60 | 64.09 | 82.33 | 94.16 | 76.40 | |
Co-SLAM | Acc. | 19.39 | 9.66 | 20.52 | 5.50 | 4.47 | 61.32 | 5.15 | 9.63 | 16.96 |
Comp. | 34.55 | 7.13 | 12.02 | 5.52 | 3.68 | 41.85 | 5.12 | 5.72 | 14.45 | |
Comp. Ratio(%) | 45.44 | 79.17 | 67.99 | 91.94 | 98.09 | 20.72 | 86.57 | 86.91 | 72.10 | |
ESLAM | Acc. | 5.44 | 6.14 | 11.16 | 4.36 | 6.07 | 6.40 | 4.05 | 5.02 | 6.08 |
Comp. | 7.43 | 9.26 | 8.89 | 6.69 | 3.38 | 12.74 | 7.63 | 10.37 | 8.30 | |
Comp. Ratio(%) | 86.68 | 83.86 | 80.29 | 91.52 | 97.19 | 72.85 | 86.33 | 83.01 | 85.22 | |
MIPS-Fusion | Acc. | 102.64 | 31.89 | 23.81 | 13.21 | 7.02 | 9.42 | 5.30 | 16.35 | 26.21 |
Comp. | 108.69 | 19.82 | 22.83 | 19.58 | 5.35 | 25.14 | 13.13 | 27.52 | 30.26 | |
Comp. Ratio(%) | 11.20 | 58.78 | 62.26 | 79.77 | 94.69 | 69.63 | 81.97 | 59.62 | 64.74 | |
BAD-SLAM | Acc. | – | 21.43 | 8.18 | – | 6.12 | 4.85 | 3.15 | – | – |
Comp. | – | 50.39 | 72.45 | – | 11.08 | 18.92 | 8.03 | – | – | |
Comp. Ratio(%) | – | 11.46 | 30.84 | – | 71.03 | 64.33 | 85.97 | – | – | |
SplaTAM | Acc. | 180.75 | 138.67 | 48.69 | 43.66 | 10.09 | 58.38 | 117.72 | 23.97 | 77.74 |
Comp. | 145.37 | 172.34 | 52.57 | 64.77 | 9.81 | 50.13 | 30.05 | 29.1 | 69.27 | |
Comp. Ratio(%) | 11.65 | 20.21 | 29.9 | 33.41 | 67.92 | 20.60 | 44.36 | 45.41 | 34.18 | |
MonoGS | Acc. | 51.60 | 60.08 | 37.56 | 28.91 | 71.95 | 74.61 | 65.42 | 48.14 | 54.78 |
Comp. | 241.12 | 360.77 | 86.99 | 64.02 | 92.75 | 124.39 | 136.52 | 67.48 | 146.76 | |
Comp. Ratio(%) | 10.86 | 12.85 | 15.82 | 29.74 | 14.73 | 9.03 | 10.81 | 27.08 | 16.37 | |
Photo-SLAM | Acc. | 12.78 | 13.10 | 17.72 | 7.67 | 7.07 | 15.77 | 16.81 | 9.76 | 12.59 |
Comp. | 15.64 | 17.80 | 23.51 | 14.08 | 8.75 | 25.15 | 15.37 | 25.24 | 18.19 | |
Comp. Ratio(%) | 58.89 | 49.20 | 39.41 | 70.81 | 75.86 | 43.57 | 59.02 | 53.79 | 56.32 | |
RTG-SLAM | Acc. | 10.73 | 10.38 | 9.32 | 7.62 | 6.64 | 10.84 | 7.77 | 10.12 | 9.18 |
Comp. | 15.10 | 10.84 | 11.57 | 9.31 | 6.07 | 13.19 | 7.94 | 12.18 | 10.78 | |
Comp. Ratio(%) | 66.29 | 65.89 | 64.06 | 78.95 | 88.35 | 64.88 | 78.79 | 70.46 | 72.21 | |
GS-ICP SLAM | Acc. | 18.77 | 12.03 | 15.14 | 9.47 | 6.46 | 12.93 | 8.86 | 9.65 | 11.66 |
Comp. | 34.46 | 12.95 | 18.70 | 20.23 | 8.75 | 30.78 | 18.49 | 21.26 | 20.70 | |
Comp. Ratio(%) | 37.19 | 63.58 | 46.75 | 65.76 | 82.87 | 52.77 | 70.21 | 61.87 | 60.13 | |
RemixFusion | Acc. | 4.88 | 6.90 | 5.93 | 3.80 | 4.77 | 5.06 | 3.59 | 3.88 | 4.85 |
Comp. | 5.39 | 4.93 | 4.94 | 4.12 | 3.58 | 5.71 | 3.68 | 3.11 | 4.43 | |
Comp. Ratio(%) | 92.13 | 95.10 | 92.94 | 94.73 | 98.26 | 91.05 | 95.48 | 98.72 | 94.80 | |
RemixFusion-lite | Acc. | 6.51 | 8.10 | 6.72 | 4.49 | 4.72 | 5.19 | 3.76 | 4.05 | 5.44 |
Comp. | 7.21 | 6.46 | 5.31 | 4.44 | 3.85 | 6.18 | 4.09 | 3.62 | 5.15 | |
Comp. Ratio(%) | 85.13 | 86.43 | 91.11 | 93.79 | 97.7 | 89.51 | 93.68 | 97.97 | 91.92 |
Methods | Replica | ScanNet | TUM RGB-D | Avg. | ||||||||||||||||||||
r0 | r1 | r2 | o0 | o1 | o2 | o3 | o4 | Avg. | FPS | 0000 | 0059 | 0106 | 0169 | 0181 | 0207 | Avg. | FPS | desk | xyz | office | Avg. | FPS | ||
iMAP∗ | 70.1 | 4.5 | 2.2 | 2.3 | 1.7 | 0.5 | 58.4 | 2.6 | 17.8 | 0.2 | 56.0 | 32.1 | 17.5 | 70.5 | 32.1 | 11.9 | 36.7 | 0.2 | 7.2 | 2.1 | 9.0 | 6.1 | 0.1 | 20.2 |
NICE-SLAM | 1.7 | 2.0 | 1.6 | 1.0 | 0.9 | 1.4 | 4.0 | 3.1 | 2.0 | 1.1 | 8.6 | 12.3 | 8.1 | 10.3 | 12.9 | 5.6 | 9.6 | 0.6 | 2.7 | 1.8 | 3.0 | 2.5 | 0.2 | 4.7 |
Co-SLAM | 0.6 | 0.9 | 1.2 | 0.5 | 0.5 | 2.0 | 1.6 | 0.7 | 1.0 | 6.6 | 7.1 | 11.1 | 9.4 | 5.9 | 11.8 | 7.1 | 8.7 | 5.2 | 2.7 | 1.9 | 2.4 | 2.3 | 4.8 | 4.0 |
ESLAM | 0.7 | 0.7 | 0.5 | 0.6 | 0.6 | 0.6 | 0.7 | 0.6 | 0.6 | 7.3 | 7.3 | 8.5 | 7.5 | 6.5 | 9.0 | 5.7 | 7.4 | 2.3 | 2.5 | 1.1 | 2.4 | 2.0 | 0.2 | 3.4 |
MIPS-Fusion | 1.1 | 1.2 | 1.1 | 0.7 | 0.8 | 1.3 | 2.2 | 1.1 | 1.2 | 2.8 | 7.9 | 10.7 | 9.7 | 9.7 | 14.2 | 7.8 | 10.0 | 3.1 | 3.0 | 1.4 | 4.6 | 3.0 | 3.0 | 4.7 |
DROID-SLAM | 0.4 | 0.4 | 0.4 | 0.3 | 0.3 | 0.5 | 0.6 | 0.5 | 0.4 | 22.6 | 8.4 | 7.8 | 9.7 | 10.8 | 10.7 | 6.5 | 9.0 | 13.9 | 2.2 | 1.4 | 1.8 | 1.8 | 32.3 | 3.7 |
Point-SLAM | 0.6 | 0.4 | 0.4 | 0.4 | 0.5 | 0.5 | 0.7 | 0.7 | 0.5 | 0.3 | 10.2 | 7.8 | 8.7 | 22.2 | 14.7 | 9.5 | 12.2 | 0.2 | 2.6 | 1.3 | 3.2 | 2.4 | 0.1 | 5.0 |
MonoGS | 0.3 | 0.2 | 0.3 | 0.4 | 0.2 | 0.3 | 0.1 | 0.8 | 0.3 | 0.7 | 9.8 | 32.1 | 8.9 | 10.7 | 21.8 | 7.9 | 15.2 | 1.7 | 1.6 | 1.4 | 1.5 | 1.5 | 1.4 | 5.7 |
SplaTAM | 0.3 | 0.4 | 0.3 | 0.5 | 0.3 | 0.3 | 0.3 | 0.6 | 0.4 | 0.3 | 12.8 | 10.1 | 17.7 | 12.1 | 11.1 | 7.5 | 11.9 | 0.5 | 3.4 | 1.2 | 5.2 | 3.3 | 0.3 | 5.2 |
LoopSplat | 0.3 | 0.2 | 0.2 | 0.2 | 0.2 | 0.5 | 0.2 | 0.3 | 0.3 | 0.5 | 6.2 | 7.1 | 7.4 | 10.6 | 8.5 | 6.6 | 7.7 | 0.6 | 2.1 | 1.6 | 3.2 | 2.3 | 0.6 | 3.4 |
Photo-SLAM | 0.5 | 0.4 | 0.3 | 0.5 | 0.4 | 1.3 | 0.8 | 0.6 | 0.6 | 28.6 | 8.3 | 6.7 | 9.2 | 8.5 | 78.8 | 7.5 | 19.8 | 22.3 | 2.6 | 0.3 | 1.0 | 1.3 | 21.3 | 7.3 |
RTG-SLAM | 0.2 | 0.2 | 0.1 | 0.2 | 0.1 | 0.2 | 0.2 | 0.2 | 0.2 | 5.9 | 125.1 | 109.1 | 128.8 | 7.8 | 28.1 | 6.9 | 67.6 | 2.9 | 1.7 | 0.4 | 1.1 | 1.1 | 3.8 | 23.0 |
GS-ICP SLAM | 0.2 | 0.2 | 0.1 | 0.2 | 0.1 | 0.2 | 0.2 | 0.2 | 0.2 | 22.4 | 78.3 | 94.5 | 41.2 | 112.6 | 59.8 | 20.1 | 67.8 | 24.3 | 2.7 | 1.8 | 2.7 | 2.4 | 24.2 | 23.5 |
RemixFusion | 0.5 | 0.4 | 0.3 | 0.2 | 0.5 | 0.5 | 0.4 | 0.4 | 0.4 | 14.3 | 6.9 | 10.3 | 9.3 | 6.7 | 15.4 | 7.6 | 9.4 | 12.0 | 2.3 | 1.8 | 2.4 | 2.2 | 10.7 | 4.0 |
Figure 8. Fully described in the text.
Figure 7 displays the reconstruction with estimated camera poses. Significant distortion is noticeable in the reconstructions produced by other methods, whereas our approach achieves accuracy and consistency. This underscores the precision of our pose estimation proposed in Section 3.3 and Section 3.4. Furthermore, the colorized trajectory in Figure 10 demonstrates the robustness and accuracy of our pose estimation even in very large-scale scenes, contrasting with failures or substantial drifts observed in alternative methods. We highly recommend readers refer to our supplementary video and materials for more comprehensive visualization and experimental results.
Table 3 shows the tracking comparison of two large-scale sequences of uHumans2. The apartment contains three floors and the office is more than , which presents significant challenges. Our method is the most accurate in terms of tracking, exceeding the SOTA method (ESLAM) by 12%, superior to the others on both sequences. DROID-SLAM, utilizing layers of powerful dense bundle adjustment, derives the second-best tracking accuracy on apartment and the third-best accuracy on office, indicating great robustness. Corresponding rendering results are ignored, since DROID-SLAM omits reconstruction. Notably, MonoGS is the best 3DGS-based method on uHumans2, with tracking errors of 6.92cm (apartment) and 26.21cm (office). The increasing error in the larger scene (office) indicates the generalization limitation for 3DGS-based methods. In contrast, our method is robust on both sequences. Moreover, our method is a real-time system, while the best implicit method (ESLAM) and 3DGS-based method (MonoGS) are hard to meet the real-time requirements.
Table 5 shows the quantitative results on room-level datasets, including Replica, ScanNet, and TUM RGB-D. Our method is the second-best on sequence scene000000 on ScanNet, which is challenging. Our method is comparable to the SOTA on average, and there is only a 0.6cm decrease in tracking accuracy compared to ESLAM. Our method performs better than ESLAM on Replica with 37.1 improvement. While ESLAM achieves the best performance on average, its system FPS is about 3, and does not meet the real-time requirements, which is important for the online SLAM system. Similar to ESLAM, LoopSplat showcases remarkable accuracy and robustness, which achieves the second-best tracking accuracy in this room-level benchmark thanks to the powerful loop closure modules. However, the modules are computationally expensive, and the efficiency is significantly limited (¡1 FPS). Our method achieves good tracking accuracy in large-scale scenes with relatively few iterations. Room-level scenarios require fewer optimizations, where we obtain comparable performance and great running efficiency (3 times faster than ESLAM and 21 times faster than LoopSplat) at the same time with the residual-based mixed representation. Therefore, our method is still superior to the others in terms of the online SLAM setting. Note that DROID-SLAM achieves higher FPS by prioritizing camera tracking over dense reconstruction. While the 3DGS-based methods like RTG-SLAM and GS-ICP SLAM are superior to other methods on Replica, they significantly fall short in real-world datasets like ScanNet and TUM RGB-D. This demonstrates the inherent instability of their camera tracking when exposed to real-world input noise. SplaTAM and MonoGS both use the rendering loss as the objective function, and demonstrate robustness in real-world room-level scenarios. Detailed comparisons of room-scale sequences on FastCaMo-Synth (noise-free), are provided in the supplementary materials.
Methods | Metrics | |||
PSNR | SSIM | LPIPS | D-L1 | |
NICE-SLAM | 20.88 | 0.940 | 0.195 | 0.173 |
Co-SLAM | 24.33 | 0.970 | 0.163 | 0.052 |
ESLAM | 22.41 | 0.943 | 0.150 | 0.145 |
MIPS-Fusion | 23.30 | 0.962 | 0.193 | 0.044 |
SplaTAM | 16.37 | 0.734 | 0.287 | 0.362 |
MonoGS | 19.93 | 0.690 | 0.468 | 1.156 |
Photo-SLAM | 23.45 | 0.962 | 0.154 | 1.085 |
RTG-SLAM | 23.81 | 0.966 | 0.163 | 0.287 |
GS-ICP SLAM | 26.23 | 0.980 | 0.118 | 0.223 |
RemixFusion | 24.65 | 0.971 | 0.154 | 0.031 |
RemixFusion can preserve the detailed rendering of not only textures but also geometry.
4.2.3. 3D Reconstruction.
The traditional explicit methods, including ElasticFusion and BundleFusion, can not succeed in finishing the tracking and reconstruction for almost all scenes. Therefore, we only report the results of the implicit methods, BAD-SLAM, and the explicit 3DGS-based method. Note that the reconstruction mesh for 3DGS-based methods (SplaTAM, MonoGS, RTG-SLAM, and GS-ICP SLAM) is obtained using TSDF Fusion with rendered RGB-D images following 2DGS (Huang et al., 2024b). Implicit methods often produce noisy meshes in empty spaces, posing challenges for fair comparisons. Following (Zhu et al., 2022; Wang et al., 2023), we use the estimated camera poses to cull the mesh with ground-truth depths for evaluation. Ground-truth meshes obtained from LiDAR scans use the same strategies with ground-truth poses. The evaluation is performed three times, and the average results are reported. Due to the scaling differences between the results of BAD-SLAM and ground-truth meshes, we align the output point clouds with ground-truth meshes before evaluation.
In Table 4, our method outperforms the SOTA implicit method by 20.2 and the SOTA 3DGS-based method by 47.2 for the reconstruction accuracy on BS3D, and there is 46.6 improvement in completeness compared to the SOTA methods. Although BAD-SLAM is the most accurate on lounge and study, its completeness suffers due to explicit surfel representations. Our method exhibits marginally lower accuracy on these two sequences (0.21cm and 0.44cm decrease, respectively) but significantly outperforms other methods on the remaining sequences. Moreover, the completeness ratio of ours is over 90 for all sequences, surpassing the other approaches by over 9.5. This further validates the effectiveness of our residual-based mixed representations. ESLAM demonstrates robust 3D reconstruction on BS3D, but is still worse than RemixFusion-lite. SplaTAM, MonoGS and Photo-SLAM exhibit poor reconstruction, primarily due to the discontinuity in depth rendering (Table 6) and the fragile tracking in large-scale scenes (Table 1 and Table 3).
Figure 7 presents the reconstruction results, with an overview in the first row and a detailed zoom-in in the second row. Our methods can preserve the finest geometric details in these challenging large-scale scenes. In contrast, alternative methods struggle with unstable pose estimation and surface distortions, highlighting the effectiveness of our residual-based mixed representations. For example, in the second row of Figure 7, the comparisons of the white lines on the floor and black chairs illustrate that our reconstruction offers superior details in both texture and geometry.
Figure 8. Fully described in the text.
Figure 10 presents additional results about the large-scale scenes, which are challenging for both tracking and reconstruction. The colorized trajectory means the scanning timeline: red denotes the start and blue denotes the end. The area of dining on BS3D exceeds and there are multiple staircases and floors. Our method successfully finished the pose estimation and detailed reconstruction, whereas other approaches failed either in tracking or returning to the starting position, resulting in reconstruction distortions. For instance, MIPS-Fusion exhibits upward leaning and significant distortion in reconstruction upon returning to the second floor, despite employing loop closure for sub-maps. Note that the input depths for the floor of this scene are missing in the beginning, leading to some empty holes in the floor. The second column in Figure 10 shows the results of building2 on FastCaMo-Large, where there are three floors. Other methods notably fail on the third floor, where the right corridor should appear flat and straight. The third column in Figure 10 is a challenging self-captured sequence with faster camera motion, posing challenges for both tracking and mapping. RemixFusion demonstrates adaptability with robust pose estimation and efficient reconstruction, outperforming all the other methods. There are severe distortions for MIPS-Fusion at the end of the trajectory (blue), and the reconstruction is of great noise due to the heavy and poorly aligned sub-maps. Our reconstruction is clean and more accurate, with FPS 4 times faster than MIPS-Fusion. The other methods can succeed in reconstructing the first half of the sequence but fail in the middle, demonstrating the importance of robustness in large scenes.
Methods | Metrics | Training Views | Novel Views | Diff. |
GS-ICP SLAM | PSNR | 29.68 | 16.85 | -12.83 |
SSIM | 0.954 | 0.817 | -0.103 | |
LPIPS | 0.074 | 0.237 | -0.163 | |
D-L1 | 0.246 | 3.002 | -2.756 | |
RemixFusion | PSNR | 28.53 | 21.14 | -7.39 |
SSIM | 0.991 | 0.963 | -0.028 | |
LPIPS | 0.093 | 0.166 | -0.073 | |
D-L1 | 0.021 | 0.381 | -0.36 |
RemixFusion is the most accurate and robust approach in large-scale scenes, where there are different and challenging situations for SLAM.
The quality comparison of office on uHumans2 is shown in Figure 9. The color of the trajectory indicates the scanning time. The reconstruction of our method is the most compact and detailed, while there is noticeable blurring in some areas or distortion. (see the zoomed-in areas marked by red rectangles). Our method is capable of preserving the texture of the floor as well as the geometric structures of fine-grained objects on the table, which is hard for other methods. The most accurate, for tracking, 3DGS-based baseline (MonoGS) is hard to successfully reconstruct the entire sequence, and there is noticeable distortion and noise in geometry.
4.2.4. 2D Rendering.
The 2D rendering comparison leverages the estimated poses from different methods to better focus on the scene reconstruction. Table 6 shows the rendering comparison of training views on BS3D. Every 10 frames of each scene in the dataset are evaluated, and pixels with missing depth are filtered. Although GS-ICP SLAM, as the representative 3DGS-based explicit method, excels in photometric metrics such as PSNR, SSIM, and LPIPS, our method achieves superior photometric rendering compared to all implicit methods and the most accurate geometric rendering on average, as illustrated by D-L1. The D-L1 values of ESLAM are inferior to those of Co-SLAM and MIPS-Fusion, demonstrating the less detailed geometric reconstruction of the utilized tri-planes. Our method significantly surpasses GS-ICP SLAM by 19.2cm in D-L1. Among the 3DGS-based methods, RTG-SLAM is the second-best method regarding PSNR and D-L1, which uses significantly fewer 3DGS, albeit at the cost of reduced reconstruction accuracy. Photo-SLAM employs a hyper-primitive map that achieves comparable RGB rendering quality to implicit methods, yet exhibits inferior depth rendering performance, indicating potential overfitting of rendering and unsatisfactory geometric reconstruction.
Methods | ATE RMSE | F1-score | FPS | GPU mem. |
NICE-SLAM | 10.4 | 54.39 | 0.5 | 9.5G |
Co-SLAM | 48.8 | 18.02 | 5 | 5.6G |
MIPS-Fusion | 11.3 | 70.70 | 3 | 7.3G |
ESLAM | 6.3 | 78.95 | 2 | 15.8G |
BAD-SLAM | 8.3 | 75.94 | 28 | 9.9G |
SplaTAM | 621.7 | 17.35 | 0.1 | 21.3G |
MonoGS | 317.2 | 7.46 | 0.8 | 15.6G |
Photo-SLAM | 24.64 | 44.42 | 21.5 | 9.1G |
LoopSplat | 149.08 | 19.21 | 0.3 | 22.0G |
RTG-SLAM | 12.0 | 66.35 | 2.1 | 10.1G |
GS-ICP SLAM | 24.0 | 55.64 | 28 | 16.2G |
RemixFusion | 4.2 | 90.25 | 12 | 9.8G |
RemixFusion-lite | 4.6 | 89.50 | 25 | 8.0G |
GS-ICP SLAM is good at 2D photometric rendering, but its geometric rendering (D-L1) is notably inferior compared to other alternatives. This can be attributed to the reason that 3DGS-based representations achieve high-fidelity rendering through the fitting of 2D training views, indicating less attention to the 3D geometry in the online setting. This can also be proved by the metrics of 3D reconstruction accuracy and completeness presented in Table 4. In terms of real-time reconstruction, 3DGS-based methods are easy to achieve high-fidelity RGB rendering. However, they struggle to get accurate geometric reconstruction (see the D-L1 metric in Table 6). For example, RTG-SLAM performs well on the scene study of BS3D with 6.7cm in ATE RMSE (shown in Table 1). However, the 3D reconstruction quality (shown in Table 4) of it is not consistent with the performance of tracking. For comparison, Co-SLAM achieves 6.5cm on the scene study of BS3D in ATE RMSE, which is similar to RTG-SLAM; the 3D reconstruction of Co-SLAM is much better than RTG-SLAM regarding the accuracy and completeness. Note that RTG-SLAM focuses on the 3DGS optimization using the tracking module from other methods; its FPS and reconstruction are still inferior to Co-SLAM. Improving geometric reconstruction quality for 3DGS-based methods often requires more optimization iterations, which is generally unacceptable for real-time SLAM.
Furthermore, the comparison of novel view synthesis in Table 7 demonstrates that our residual-based representation significantly outperforms 3DGS in photometric rendering (PSNR) by 25 and surpasses it in geometric rendering (D-L1) by over 7 times. This indicates that the 3DGS-based methods are less geometry-aware in the online setting. Note that the sequence cafeteria-nvs of BS3D in Table 7 uses the manually partitioned train (50) and test images (50), with ground-truth poses excluding the effects of tracking. Overall, our residual-based representations surpass both implicit methods and explicit methods, including 3DGS, taking efficiency and performance into consideration in large scenes.
Figure 8 illustrates the 2D rendering results of our method and GS-ICP SLAM (the second-best method on BS3D), which indicates that our method is more stable in photometric rendering and much better in geometric rendering than the 3DGS-based method. For example, our rendering of the small table near the green number 10 on foobar is more detailed and complete. We attribute this to the residual-based representation of RemixFusion and the direct supervision of 3D space. The novel view synthesis comparison is presented at the bottom of Figure 8. Our method remains accurate, while there is notable noise for GS-ICP SLAM, which proves that GS-ICP SLAM only overfits the 2D training views without accurately reconstructing the real 3D geometry.
Table 3 shows the rendering comparison on uHumans2 dataset. Our method outperforms the SOTA in terms of tracking by 12 and D-L1 by 12.3. Our method is the best in terms of PSNR on office, surpassing the SOTA by 0.33dB. While ESLAM is the second-best method for D-L1, its system FPS is over three times slower than our method. For 3DGS-based methods, MonoGS demonstrates the best RGB rendering on apartment (surpassing the second-best one by 4.49 for PSNR), but its depth rendering is far worse than other methods. This also indicates that 3DGS-based methods are less geometry-aware in the online setting, as mentioned above. Overall, our method is the best real-time method in terms of tracking and rendering on average.
RemixFusion can reconstruct the whole scene within extremely few optimization iterations.
4.2.5. Runtime and Memory Analysis
Table 8 illustrates the system FPS and the performance of both tracking and mapping. To ensure a comprehensive comparison, we focus on the scene lounge from BS3D, which is successfully handled by most methods. Note that the FPS of explicit methods like ElasticFusion and BundleFusion is high, but both of them fail in this scene. Therefore, they are not compared in this table. Our method achieves the best trade-off between GPU memory usage and accuracy. Note that our method maintains relatively consistent GPU memory usage regardless of scene scale, whereas explicit methods increase GPU memory usage with additional input frames in larger scenes. Thanks to the residual-based mapping, our method demonstrates superior system FPS compared to implicit methods, and the lightweight version of our method can run at 25 FPS with GB GPU memory usage, which is as fast as the explicit alternatives and requires less computational overhead.
Conversely, approaches like Co-SLAM and MIPS-Fusion employ hash schemes or multiple sub-maps to reduce memory consumption but struggle with efficiency and robustness in memorizing large-scale scenes. Methods like ESLAM, SplaTAM, MonoGS are not memory-friendly, requiring half-resolution images as input. Additionally, their system FPS is significantly lower than that of other methods. Similarly, LoopSplat is not memory-efficient in large scenes and exhibits running inefficiency (¡1 FPS), primarily due to the computationally expensive modules of loop closure and post-refinement. Photo-SLAM, leveraging ORB-SLAM3 for tracking and a map with hyper primitives for mapping, is lightweight and efficient. However, Photo-SLAM indicates worse tracking and reconstruction accuracy on large scenes. The traditional explicit method, like BAD-SLAM, is real-time but lacks robustness. GS-ICP SLAM, as the representative of 3DGS-based methods, is fast but compromises memory efficiency. Although RTG-SLAM leverages efficient techniques to reduce the redundant 3DGS and requires less GPU memory in the process, it requires large GPU memory for global optimization at the end of tracking. Our method stands out as the most efficient system, coupled with the most accurate pose estimation as well as reconstruction, considering both FPS and GPU memory usage.
4.3. Ablation Studies
We perform ablation studies mainly on the BS3D (Mustaniemi et al., 2023) dataset and report the average results of all scenes to evaluate the effectiveness of the modules proposed in RemixFusion. We analyze the necessity of the key components proposed in our methods for validation, which are residual-based mapping, residual-based bundle adjustment, and gradient amplification, respectively.
4.3.1. Residual-based Mapping
One key innovation lies in the residual-based mapping, crucial for enabling real-time dense reconstruction of large-scale scenes in our system. To validate this design, we first compare our mixed method against using solely the implicit module without residuals. Parameters are the same except for the residual-based representations. Moreover, we demonstrate how the reconstruction quality varies according to the desired mapping FPS.
With the proposed gradient amplification, the bundle adjustment is encouraged to explore more and find more global optimal solutions.
RBA | TBA | GA | ATE RMSE |
6.39 | |||
✓ | 5.68 | ||
✓ | 5.74 | ||
✓ | ✓ | 5.14 | |
✓ | ✓ | 4.61 |
We evaluate 8 sequences of BS3D and report the average D-L1 values in Figure 11. For a fair comparison focused solely on reconstruction, excluding tracking, we use the ground-truth poses for reconstruction, and no bundle adjustment for poses is utilized. As illustrated by Figure 11, the D-L1 values (cm) of our method with the residual design exhibit slower degradation even with higher required FPS. This indicates that our mapping module with residuals achieves faster convergence, and the D-L1 values are below 5cm even if the mapping process is running at 86 FPS. Moreover, the result of the mixed representations is significantly better than that of the coarse map, indicating the effective residual-based learning based on the coarse map. Figure 11 also shows the reconstruction comparison when the mapping FPS is 86, which means only about iterations for mapping are performed for about 2000 frames. This is quite challenging, yet the results demonstrate that the residual-based mapping module can achieve meaningful reconstruction details with minimal optimization. The experiments prove that the proposed residual-based mapping is efficient in preserving the detailed reconstruction within a limited time.
4.3.2. Residual-based bundle adjustment
Another key insight of our method is the residual-based bundle adjustment (RBA for short). Here we evaluate the effectiveness of our residual-based bundle adjustment for pose refinement. We compare our method to the traditional bundle adjustment that optimizes the individual pose variables (TBA for short). Additionally, we validate the effectiveness of our proposed gradient amplification (GA for short).
Compared to the commonly used bundle adjustment, which individually optimizes the pose of each frame, our proposed method employs a single MLP for 6D residual pose prediction, offering enhanced global consistency. Table 9 illustrates that the average ATE RMSE without bundle adjustment (baseline) is 6.39cm. Applying the RBA instead of TBA for pose refinement reduces the error marginally from 5.74cm to 5.68cm. Introducing GA to help bundle adjustment results in improved pose estimation for both TBA and RBA. Moreover, using RBA with GA reduces the error from 5.68cm to 4.61cm, resulting in an improvement of over 1cm. This proves the effectiveness of the proposed gradient amplification, which is simple yet necessary for the residual-based BA in large-scale scenes. Figure 14 demonstrates the comparison of different methods on foobar of BS3D, which validates the effectiveness and necessity of the proposed residual-based bundle adjustment and gradient amplification. In summary, the proposed residual-based BA can focus on detailed refinement of the initial camera poses and significantly improve the accuracy of pose estimation in large-scale scenes.
The threshold of 2 for gradient amplification improves all the sequences on BS3D, and the average improved performance of it is the highest.
The proposed residual-based bundle adjustment coupled with gradient amplification improves the accuracy and robustness of camera tracking in large-scale scenes.
4.3.3. Adaptive Gradient Amplification
A key design in the residual-based BA is the proposed adaptive gradient amplification used in BA. As illustrated in Eq. 17, this technique amplifies the optimization gradients derived from the reconstructed surface, allowing the BA process to circumvent local minima and thus achieve more globally consistent solutions in real-world large-scale scenarios. The visualization of the optimization is shown in Figure 12. This design aims to mitigate the risk of falling into local optima, a common challenge in large-scale scenes. Visualization of the optimization of BA on waiting and corridor on BS3D is shown in Figure 12, where the 6D poses of all frames are visualized in 2D and colorized by the cost functions. Without the proposed gradient amplification, the optimization tends to get trapped in the local minimum (illustrated by the blue dashed line). However, the proposed gradient amplification enables BA to escape from the local minimum and obtain a better solution (orange dashed line).
Furthermore, we compare our method (GA) with the manually designed amplification (MA) and the randomized amplification (RA). Figure 13 presents bundle adjustment (BA) results under various truncation thresholds compared to the baseline (without GA). The positive improvement indicates the effectiveness of gradient amplification. Results of eight scenes of BS3D are reported as the range of the bar chart. GA achieves the best results on average with , showing positive improvements across all scenes. MA-5 and RA-5 indicate that the camera poses are guided towards the directions with 5cm, where the cameras are facing or randomized directions. Our proposed GA yields the most accurate pose estimation, while MA also improves the performance, which is less than GA. RA fails to improve the performance of bundle adjustment for all sequences. In summary, the improved results prove the effectiveness of the proposed gradient amplification.
Figure 15. Fully described in the text.
4.4. Limitation and Failure Cases.
There is a primary limitation of our method. Robust pose estimation becomes challenging when a significant portion of the observed depth images is missing because our approach heavily depends on depth information. Consequently, our method can not be applied in scenarios where only RGB data is available. Failure cases of our method are shown in Figure 15. The majority of observed depth information in the middle of the sequence (highlighted by red rectangles) is missing, making it hard and unstable for both pose estimation and reconstruction.
5. CONCLUSION
It is still a challenging problem to perform online dense reconstruction for a large-scale environment with fine-grained geometry details preserved. It is critical to formulate a memory-friendly scene representation that can support efficient and high-quality tracking and mapping. With our work, we wish to bring to the community’s attention that a residual-based mixture is a proper way to take advantage of both explicit and implicit formulations. By reducing the learning burden on implicit networks through coarse-grained explicit storage, we have significantly accelerated the efficiency of online reconstruction. This enhancement allows our residual-based framework to preserve more reconstruction details while ensuring real-time performance. Note that the residual idea also inspires a new approach to pose estimation, where we optimize only the pose changes during multi-frame joint optimization, thus reducing the network’s learning complexity. The evaluation comparison between our method and other alternatives demonstrates the superiority of RemixFusion in tracking accuracy and mapping quality for large-scale scene reconstruction. While our method can handle large-scale scene reconstruction with limited memory cost, it is still worth doing as future work to make this mixed representation dynamically scalable, which may be able to support larger online dense reconstruction at a city block level.